OpenStudy (wyattp17):

Graph the circle with center at (4, 0) and radius 3. Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (wyattp17):

@Elsa213

Elsa213 (elsa213):

I just answered this question....

Elsa213 (elsa213):

the equation of a circle is (x-a)^2+(y-b)^2=r^2 r is the radius and (a,b) the center of the circle

Elsa213 (elsa213):

the equation is (x-4)^2+(y-0)^2=3^2 or x^2+y^2-8x=-7

Elsa213 (elsa213):

the graph is easy, you know the center and the radius, to find the four points you just pick a value of x and solve for y

Elsa213 (elsa213):

http://prntscr.com/4b6c3h

Elsa213 (elsa213):

pick values in the interval [1,7] because of the diameter

Elsa213 (elsa213):

I got this from: http://openstudy.com/study#/updates/5301254ce4b018b58423fe2e

OpenStudy (wyattp17):

@Elsa213 what are 4 points i got two